How much of an 8% saline solution should be added to 600ml of a 3% solution to produce a 5% solution?
Answer by stanbon(75887)

How much of an 8% saline solution should be added to 600ml of a 3% solution to produce a 5% solution?
---------------
0.08x + 0.03*600 = 0.05(x+600)
-----
Multiply thru by 100 to get:
8x + 3*600 = 5x + 5*600
---
3x = 2*600
---
x = 400ml (amt of 8% solution needed)
===========================================
Cheers,
Stan H.